[Brahms-l] Shift report 20040207 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sat Feb 07 2004 - 15:59:45 EST
------------------------------------------------------------
Date: Sat Feb  7 15:59:44 2004, Shift: 08:00-16:00
Supervisor: rk, with: 
Log book#15 pages 50 - 50

Shift Summary:
------------------------------------------------------------
1. Beam for less than 1 hour.
2. Pedestals 10196,7. Data run 10198 with FS@2.3@1/sqrt2B and MRS@45@12A.
3. Had some problems with TPM1, during turning it on, it tripped. After restarting brahvo, it came back to normal.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 7 16:00:13 2004

This archive was generated by hypermail 2.1.8 : Sat Feb 07 2004 - 16:00:29 EST